# Quantum simulation of strong Charge-Parity violation and Peccei-Quinn mechanism

### v1

@@ I End Matter
Four sites and three links (7 qubits) model. We consider a four-site open chain with three gauge links. Matter (staggered) sites are $x=0,1,2,3$ and links are $\ell=4,5,6$ connecting (x,x+1)(x,x{+}1). In a spin-$\tfrac{1}{2}$ quantum link model,
@@
$H=H_{m}+H_{\rm hop}+H_{\theta}+H_{G}+\text{const.}$ (40)
Mass term.
With the standard staggered sign
@@
Gauge electric energy with $\theta$-term.
Using the shifted quadratic form
$H_{\theta}=\frac{g^{2}}{2}\sum_{\ell=4}^{6}\!\left(E_{\ell}-\frac{\theta}{2\pi}\right)^{2},\qquad E_{\ell}=\frac{I-Z_{q_{\ell}}}{2}.$ (42)
Gauge-invariant hopping.
With JW strings $\psi_{x}=(\prod_{y<x}Z_{q_{y}})\,\sigma^{-}_{q_{x}}$ and link $U_{\ell=x+4}$ on (x,x+1)(x,x{+}1), the cancellations yield
@@
$=\frac{w}{8}\!\sum_{x=0}^{2}\!\Big(X_{q_{x}}X_{q_{x+1}}X_{q_{\ell}}+Y_{q_{x}}Y_{q_{x+1}}X_{q_{\ell}}$
$\hskip 28.45274pt-X_{q_{x}}Y_{q_{x+1}}Y_{q_{\ell}}-Y_{q_{x}}X_{q_{x+1}}Y_{q_{\ell}}\Big).$ (43)
Gauss’s law and penalty.
For open boundaries,

### v2

@@ Appendix B Appendix B: Ground-state preparation via FALQON
Appendix C Appendix C: Four sites and three links (7 qubits) model
We consider a four-site open chain with three gauge links. Matter (staggered) sites are $x=0,1,2,3$ and links are $\ell=4,5,6$ connecting (x,x+1)(x,x{+}1). In a spin-$\tfrac{1}{2}$ quantum link model,
@@ Appendix C Appendix C: Four sites and three links (7 qubits) model
$H=H_{m}+H_{\rm hop}+H_{\bar{\theta}}+H_{G}+\text{const.}$ (41)
First, the mass term with the standard staggered sign
@@
For the Gauge electric energy with $\bar{\theta}$-term, using the shifted quadratic form
$H_{\theta}=\frac{g^{2}}{2}\sum_{\ell=4}^{6}\!\left(E_{\ell}-\frac{\bar{\theta}}{2\pi}\right)^{2},\qquad E_{\ell}=\frac{I-Z_{q_{\ell}}}{2}.$ (43)
For the Gauge-invariant hopping, with JW strings, we have $\psi_{x}=(\prod_{y<x}Z_{q_{y}})\,\sigma^{-}_{q_{x}}$ and link $U_{\ell=x+4}$ on (x,x+1)(x,x{+}1), the cancellations yield
@@
$=\frac{w}{8}\sum_{x=0}^{2}\!\Big(X_{q_{x}}X_{q_{x+1}}X_{q_{\ell}}+Y_{q_{x}}Y_{q_{x+1}}X_{q_{\ell}}$-
$\hskip 56.9055ptX_{q_{x}}Y_{q_{x+1}}Y_{q_{\ell}}-Y_{q_{x}}X_{q_{x+1}}Y_{q_{\ell}}\Big).$ (44)
For Gauss’s law and penalty, with open boundaries,